Let A(3,2,1), B(5,0,2), C(3,3,0) and D(1,-6,8) be four points in R3. use vector methods to solve the following.

a) Find the smaller angle between the side AB and AC.
b) Find the shortest distance between D and ABC plane, without finding the equation of the plane
1
Expert's answer
2020-03-10T13:00:40-0400

"A(3,2,1), B(5,0,2), C(3,3,0), D(1,-6,8)"

1) Form vectors

"\overrightarrow{AB}=(5-3,0-2,2-1)=(2,-2,1)\\
\overrightarrow{AC}=(3-3,3-2,0-1)=(0,1,-1)"

We will find a scalar product

"\overrightarrow{AB}\cdot\overrightarrow{AC}=|\overrightarrow{AB}|\cdot
|\overrightarrow{AC}|\cdot cos\alpha=\\
=x_{\overrightarrow{AB}}\cdot x_{\overrightarrow{AC}}+y_{\overrightarrow{AB}}\cdot
y_{\overrightarrow{AC}}+z_{\overrightarrow{AB}}\cdot z_{\overrightarrow{AC}}"

"|\overrightarrow{AB}|=\sqrt{2^2+(-2)^2+1^2}=\sqrt{4+4+1}=3\\
|\overrightarrow{AC}|=\sqrt{0^2+1^2+(-1)^2}=\sqrt{0+1+1}=\sqrt{2}\\
\overrightarrow{AB}\cdot\overrightarrow{AC}=2\cdot0+(-2)\cdot1+1\cdot(-1)=-3\\
\overrightarrow{AB}\cdot\overrightarrow{AC}=|\overrightarrow{AB}|\cdot
|\overrightarrow{AC}|\cdot cos\alpha\\
-3=3\cdot\sqrt{2}\cdot cos\alpha\\
cos\alpha=-\frac{1}{\sqrt2}\\
\alpha=135^0"

The smaller angle between the side AB and AC

"180^0-\alpha=180^0-135^0=45^0"

2) Form vectors

"\overrightarrow{AB}=(2,-2,1)\\
\overrightarrow{AC}=(0,1,-1)\\
\overrightarrow{AD}=(1-3,-6-2,8-1)=(-2,-8,7)"

Find the volume of the pyramid "DABC"

"V=\frac {1}{3}S\cdot H\\
V=\frac{1}{6}|\overrightarrow{AB}\cdot\overrightarrow{AC}\cdot\overrightarrow{AD}|"

where "\overrightarrow{AB}\cdot\overrightarrow{AC}\cdot\overrightarrow{AD}" is the scalar triple product,

"S=\frac{1}{2}|\overrightarrow{AB}\times\overrightarrow{AC}|"

where "\overrightarrow{AC}\times\overrightarrow{AC}" is the cross product

"\overrightarrow{AB}\cdot\overrightarrow{AC}\cdot\overrightarrow{AD}=\begin{vmatrix}
 2 & -2&1 \\
 0 &1&-1\\
-2&-8&7
\end{vmatrix}=\\
=2\cdot1\cdot7+(-2)\cdot(-1)\cdot(-2)+0\cdot(-8)\cdot1-\\
-1\cdot1\cdot(-2)-(-2)\cdot0\cdot7-(-8)\cdot(-1)\cdot2=\\
=14-4+0+2-0-16=-4"

"V=\frac{1}{6}|-4|=\frac{2}{3}"

"\overrightarrow{AB}\times\overrightarrow{AC}=(\begin{vmatrix}
 -2 & 1 \\
 1 & -1
\end{vmatrix},-\begin{vmatrix}
 2 & 1 \\
 0 & -1
\end{vmatrix},\begin{vmatrix}
 2 & -2 \\
 0 & 1
\end{vmatrix})=\\
=(2-1,-(-2-0),2-0)=(1,2,2)\\
|\overrightarrow{AB}\times\overrightarrow{AC}|=\sqrt{1^2+2^2+2^2}=\sqrt9=3"

"S=\frac{1}{2}\cdot3=\frac{3}{2}\\"

"V=\frac {1}{3}S\cdot H\\
\frac{2}{3}=\frac{1}{3}\cdot\frac{3}{2}\cdot H\\
H=\frac{4}{3}"

The shortest distance between D and ABC plane is "\frac{4}{3}"